1.Given the Keynesian model:

Y  =   C + I + G + X – M

Where C = 30 + 0.8Yd; Yd = Y – T;  I = 60; T = 50; G = 50

X = 50 – 0.05Y; X is net export: 

a.Find the equilibrium national income.

b.Find the net exports balance at the equilibrium national level of income.

C.What happens to equilibrium national income and the net exports balance when net investment increases from 60 to 70?

d.Explain what happens to equilibrium national income and net exports balance when net exports function changes to 40 – 0.05Y. (5 marks each)


1
Expert's answer
2021-08-23T13:19:11-0400
"Y=30+0.8(Y-50)+60+50+50-0.05 Y"

"0.25Y=150"

a.Find the equilibrium national income.

"Y=600"

b.Find the net exports balance at the equilibrium national level of income.


"50-0.05Y=0"

"Y=1000"

C.What happens to equilibrium national income and the net exports balance when net investment increases from 60 to 70?


"50-0.05Y=60"

"Y=200"

"50-0.05Y=70"

"Y=400"

d.Explain what happens to equilibrium national income and net exports balance when net exports function changes to 40 – 0.05Y.


"Y=30+0.8(Y-50)+60+50+40-0.05Y"

"0.25Y=140"

"Y=560"